Find class boundaries, midpoint, and width for the class 18.6-21.0


1
Expert's answer
2021-09-21T11:25:54-0400

Solution:

Given the class 18.6-21.0

Upper class boundary = 21.0

Lower class boundary = 18.6

Midpoint = "\dfrac{18.6+21.0}2=19.8"

Width = 21.0 - 18.6 = 2.4